OpenStudy (anonymous):

Consider the function: F(x) = (x+4)/(x+2) Evaluate the difference quotient for the given function. Simplify your answer? F(x) - f(2)/(x-2)

OpenStudy (anonymous):

F(2)=6/4=3/2

OpenStudy (anonymous):

2(x+4)/2(x+2)-3/2(x-2) = (2x+8-3)/2(x-2) = (2x+5)/2(x-2)

OpenStudy (anonymous):

thats not right when i submitted it?

OpenStudy (anonymous):

pellet i see the mistake

OpenStudy (anonymous):

\[\frac{x+4}{x+2}-\frac{3}{2(x-2)}=\frac{(x+4)*2*(x-2)-3(x+2)}{-2(x+2)(x-2)}\]